The Hampton Inn in Stow, Ohio, has two sets of employees, the front desk staff and the housekeeping staff. There is always one front desk clerk on duty at all times, 24/7, and all of the front desk clerks make $8.00 an hour.

The housekeeping staff is also paid $8.00 an hour, but they are paid for the number of hours that it takes them to clean the rooms in the hotel. Each room takes 30 minutes to clean, but if a room has not been used the night before, it does not have to be cleaned, so the housekeeping staff work fewer hours on days when the hotel is not as full.

The hotel calculates its profitability on a per-room basis, since they have some costs that vary with the number of rooms that are occupied on a particular night. Right now, their costs other than staff wages are marketing expenses of $5,400 per month, management salaries of $60,000 per month, and general overhead (costs like utilities and insurance) of $30,000 per month.

The hotel has 50 rooms, all of which rent for the same price per night.

  1. Assuming a 30-day month, what are the hotel’s total fixed costs per day? (Be careful, read all of the cost information carefully and think before you answer this question.)
  1. What is the hotel’s variable cost per room?
  1. If the hotel fills all of its rooms for the night, what price should they charge to break even?
  1. What price should they charge to break even with only 70% of the hotel’s rooms sold for the night?
  1. What price should they charge to earn a profit margin of 25% when 70% of the hotel’s rooms are sold for the night?
Answer #1

All amounts are in $

The information about fixed and variable costs are given

1. Fixed Costs per day

Total fixed Costs

Front desk clerk salaries = 24 hrs x 8 = 192

Marketing expenses = 5,400/ 30 days = 180

Management salaries = 60,000/30 days = 2,000

General Overhead = 30,000/30 days = 1,000

Total fixed costs per day = 3,372 per day

2. Variable Cost per room per day = 4 per room per day

Housekeeping staff charges $8 per hour and it takes half an hour to clean a room. So it costs $4 to clean a room

Note : This is the only variable cost given in the question

3. Total rooms available are 50 rooms

Total fixed cost = 3,372

This must be the contribution to be obtained from the room rental service

Contribution to be obtained from a single room = 3,372/50

= 67.44

So, Price to be charged per room = Contribution to be obtained + Variable cost per unit

= 67.44 + 4

= 71.44

4. If only 70% rooms are occupied

Rooms occupied = 35 rooms

Contribution to be obtained from single room = 3,372/35

= 96.34

Price to be charged per room = 96.34 + 4 = 100.34

5. If we want to earn a Profit margin of 25% and only 70% is room occupancy

Total Cost = 3,372 + 4 x 35 rooms = 3,512

Profit margin required = 25%

So % on cost will be 33.33%

Profit = 33.33% of 3,512 = 1,171

Sales to be generated = 1,171 + 3,512 = 4,683

Price per room = 4,683/35 rooms = 133.8

Note :

Front desk clerks are 4 in numbers but each will work for 6 hours a day making it 24 hours each day at 8 per hour as cost